if a:b = 3:2, then find the value of (2a+3b):(2a+5b).
Found 2 solutions by Theo, greenestamps:
Answer by Theo(13342)

a:b can also be written as a/b.
therefore a:b = 3:2 can be written as a/b = 3/2
solve for a to get a = 3b/2
multiply both sides of this equation by 2 to get:
2a = 3b
in the ratio of (2a+3b)/(2a+5b), replace 2a with 3b to get:
(3b+3b)/(3b+5b) = 6b/8b = 6/8 = 3/4
that's the value of (2a+3b)/(2a+5b)
that's your solution.

Answer by greenestamps(13200)



Given a:b = 3:2, let a=3x and b=2x.

That is one standard way of starting working on a problem where a ratio is given like that.

Then